# Nemotron-H: A Family of Accurate and Efficient Hybrid Mamba-Transformer Models

### v2

@@ Abstract
Abstract. As inference-time scaling becomes critical for enhanced reasoning capabilities, it is increasingly becoming important to build models that are efficient to infer. We introduce Nemotron-H, a family of 8B and 56B/47B hybrid Mamba-Transformer models designed to reduce inference cost for a given accuracy level. To achieve this goal, we replace the majority of self-attention layers in the common Transformer model architecture with Mamba layers that perform constant computation and require constant memory per generated token. We show that Nemotron-H models offer either better or on-par accuracy compared to other similarly-sized state-of-the-art open-sourced Transformer models (e.g., Qwen-2.5-7B/72B and Llama-3.1-8B/70B), while being up to 3$\times$ faster at inference. To further increase inference speed and reduce the memory required at inference time, we created Nemotron-H-47B-Base from the 56B model using a new compression via pruning and distillation technique called MiniPuzzle. Nemotron-H-47B-Base achieves similar accuracy to the 56B model, but is 20% faster to infer. In addition, we introduce an FP8-based training recipe and show that it can achieve on par results with BF16-based training. This recipe is used to train the 56B model. All Nemotron-H models will be released, with support in Hugging Face, NeMo, and Megatron-LM.
@@ 1 Introduction
Overall, the Nemotron-H family of models demonstrates that hybrid models can be state-of-the-art in terms of capabilities while offering improved inference speed. Moreover, we believe FP8 pre-training and compression techniques like MiniPuzzle can make it cheaper and more efficient to create such models. To enable further research, we are releasing the models discussed in this report.
@@ 5 Alignment and Long Context
We use the Nemotron-H-8B-Base model as the starting point for this study and conduct instruction-tuning and long-context extension using a pipeline consisting of three stages. In the first stage (stage1), we perform supervised fine-tuning (SFT) over a data blend consisting of 6 million code- and math-related prompts and responses, as well as other data with general instruction-following tasks. This data blend is enhanced with long-context dependencies by sampling a small set of (prompt, response) pairs and synthetically extending them to create longer contexts. Concretely, given a (prompt, response) pair from the stage1 dataset, we extend the sample to (prompt, response, extended_prompt, extended_response) where extended_prompt and extended_response are generated using an instruction-tuned model. We repeat this process until the sample reaches the desired length (e.g., 32k tokens) and include these additional long-context examples in the stage1 dataset for SFT training.
In the second stage (stage2), we start with preference tuning on general-domain prompts instead of SFT. For preference optimization, we employ offline RPO (Sun et al., 2025). Following this step, we perform a round of RPO focused on narrow instruction following, similar to the style of prompts and responses in the IFEval Benchmark (Zhou et al., 2023).
@@
To support long contexts, we extend our stage1 SFT dataset to 512k tokens and (in similar fashion) our off-policy RPO dataset to 32k tokens. We observed that training on longer samples (e.g., 512k tokens) improved RULER (Hsieh et al., 2024) scores on shorter sequence lengths as well (e.g., 128k).
@@ 6 Conclusions
We will be releasing the Nemotron-H family of models described in this paper (along with instruct and reasoning models based off of Nemotron-H-47B-Base) with support in Hugging Face, NeMo, and Megatron-LM to facilitate further research.

### v3

@@ Abstract
Abstract. As inference-time scaling becomes critical for enhanced reasoning capabilities, it is increasingly becoming important to build models that are efficient to infer. We introduce Nemotron-H, a family of 8B and 56B/47B hybrid Mamba-Transformer models designed to reduce inference cost for a given accuracy level. To achieve this goal, we replace the majority of self-attention layers in the common Transformer model architecture with Mamba layers that perform constant computation and require constant memory per generated token. We show that Nemotron-H models offer either better or on-par accuracy compared to other similarly-sized state-of-the-art open-sourced Transformer models (e.g., Qwen-2.5-7B/72B and Llama-3.1-8B/70B), while being up to 3$\times$ faster at inference. To further increase inference speed and reduce the memory required at inference time, we created Nemotron-H-47B-Base from the 56B model using a new compression via pruning and distillation technique called MiniPuzzle. Nemotron-H-47B-Base achieves similar accuracy to the 56B model, but is 20% faster to infer. In addition, we introduce an FP8-based training recipe and show that it can achieve on par results with BF16-based training. This recipe is used to train the 56B model. We are releasing Nemotron-H base model checkpoints with support in Hugging Face and NeMo.
@@ 1 Introduction
Overall, the Nemotron-H family of models demonstrates that hybrid models can be state-of-the-art in terms of capabilities while offering improved inference speed. Moreover, we believe FP8 pre-training and compression techniques like MiniPuzzle make it cheaper and more efficient to create such models.
To enable further research, we have released the following base model checkpoints in Hugging Face and NeMo formats on the Hugging Face and NGC model repositories:
@@ 5 Alignment and Long Context
We use the Nemotron-H-8B-Base model as the starting point for this study and conduct instruction-tuning and long-context extension using a pipeline consisting of three stages. In the first stage (stage1), we perform supervised fine-tuning (SFT) over a data blend consisting of 6 million code- and math-related prompts and responses, as well as other data with general instruction-following tasks. In order to improve long context performance, part of the training in this stage is done on conversations extended to 512k tokens by concatenating shorter samples. In addition, we create long-context dependencies by selecting some existing turns in the conversation and generating new turns that reference them. These referencing turns are placed randomly throughout the extended conversations. We also generate some amount of inter-related segments connected by specific topics which are placed throughout the sample. We observed that training on these longer samples (e.g. 512k tokens) improved RULER (Hsieh et al., 2024) scores on shorter sequence lengths as well (e.g. 128k).
In the second stage (stage2), we start with preference tuning on general-domain prompts instead of SFT. For preference optimization, we employ offline RPO (Sun et al., 2025). Following this step, we perform a round of RPO focused on narrow instruction following, similar to the style of prompts and responses in the IFEval Benchmark (Zhou et al., 2023). Samples in this stage are also randomly extended up to 32k tokens using the conversations from stage1, which helps retain long-context performance during this stage.
@@ 6 Conclusions
We have released the Nemotron-H base checkpoints described in this paper with support in Hugging Face and NeMo to facilitate further research: